Parth (parthkohli):

If \(f(x) \le g(x)\) for all \(x\) on \([a,b]\) (except possibly at \(x = c\)) and \(a \le c \le b\) then \[ \lim_{x \to c} f(x) \le \lim_{x \to c} g(x)\] May I know what this means?

Parth (parthkohli):

Is this as simple as something like 'if \(f(x)\) is less than or equal to \(g(x)\), then we have \(\lim \limits_{x \to c} f(x) \le \lim \limits_{x \to c} g(x)\)'?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you want a proof for this or what ?!

Parth (parthkohli):

No, I don't. I just saw this on Paul's Notes and I got a little confused as to what it actually was supposed to mean.

OpenStudy (anonymous):

not much to this, don't think too hard about it

Parth (parthkohli):

What's that \([a,b]\) and \(c\) thing? Are [a,b] the coordinates?

OpenStudy (anonymous):

[a,b] is the set of the real numbers between a and b including a and b !

Parth (parthkohli):

Oh! That's the interval notation!

Parth (parthkohli):

Wait. How can we have \(x \in [a,b], \quad x \ne c, \quad a \le b \le c\) at the same time?

Parth (parthkohli):

Since \(x\) fulfils the value of numbers between \(a\) and \(b\); and \(c\) is between a and b, then \(c\) also is a solution!

Parth (parthkohli):

Oh. I get it now :)

Parth (parthkohli):

\(x \in [a, c) \cup (c,b]\).

OpenStudy (anonymous):

yeaah !